1. A person walks from point (2, 2, 3) to point (3,-3,1) and from there they walk to (1, 1, 1). Assuming they walk along straight lines how far did they travel? Sketch their path. 2. Plot the discrete trajectory: 1 -3 t 0s 40 t-1s 3. 4. t-2s 7. 2 4s 7 2.6t-5s
Answer #1

The distance between (2,2,3) and (3,-3,1) is find out by using,

AZ V(2-3)2 + (2 + 3)2 + (3-1)2

AB = \sqrt{ 1 + 25 + 4}

AB = \sqrt{30}

Now, the distance between (3,-3,1) and (1,1,1) is

BC = \sqrt { (1-3)^2 + (1+3)^2 + (1-1)^2}

BC = \sqrt { 4 + 16 + 0}

BC = \sqrt {20}

Hence, the total distance is

D = \sqrt {20} + \sqrt {30}